OpenStudy (anonymous):

Dorothy is 14 years younger then Rita. Ten years ago, rita was 3 times as old as dorothy was then. how old is each now? step by step plz :)

OpenStudy (anonymous):

I think I've got the answer. See what you think. We have that D (Dorothy) = R (Rita) - 14 and we are told that 10 years ago Rita's age was equal to 3 times Dorothy's, so (R-10) = 3(D-10) R-10 = 3D - 30 R = 3D - 20 Now we can find R because we know that D = R-14 and by substituting this into the equation we have only one variable. R = 3(R-14) - 20 R = 3R - 42 - 20 -2R = -62 therefore R = 31 You can use this to find D and then verify that 10 years ago R was equal to 3D

OpenStudy (anonymous):

thank u